OpenStudy (anonymous):

How do you find the substitution for these? 2x + y = 7 y - x = 1

OpenStudy (anonymous):

solve the second equation for y and then plug into the first one

OpenStudy (andriod09):

do yo know how to add equations yet?

OpenStudy (anonymous):

Solve The Second Equation For Y And Then Plug Into The First One

OpenStudy (anonymous):

I kind of know how to add them, but the set up of it being in two is confusing.

OpenStudy (andriod09):

y-x=1 find y in that equation. Then plug y into the top equation and then find x.

OpenStudy (anonymous):

basically when doing substitution you want to solve one equation for one variable (either x or y in this case) and then plug it into the equation you did not use

OpenStudy (anonymous):

Y-x=1 Find Y In That Equation. Then Plug Y Into The Top Equation And Then Find X

OpenStudy (anonymous):

Why is damn just restated things people have already said? lol

OpenStudy (anonymous):

well before I continue on, the question given is : Which of the following would not be the result of a substitution in the following system? these are the options: 7 - 2x - x = 1 2x + x + 1 = 7 2(y + 1) + y = 7

OpenStudy (anonymous):

restating*

OpenStudy (anonymous):

try doing the suggested manipulation stated above and then see what you get..

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Ill work it out for you... (1) 2x+y=7 (2) y-x=1 Solving (2) for y yields: y=x+1--- we can call this (3) now plugging (3) into (1) : 2x+(x+1)=7

OpenStudy (anonymous):

2x+4=7?? and you didn't have to work it out, but thank you for doing so.

OpenStudy (anonymous):

Ooops, I misread, that was the answer at the bottom. Thank you for the help, I now know how to set it.